OpenStudy (anonymous):

Consider the quadratic function f(x)=1/2x^2+2x-6. express f in vertex form find the extreme value and classify it as a local maximum or local minimum. Consider the quadratic function f(x)=1/2x^2+2x-6. express f in vertex form find the extreme value and classify it as a local maximum or local minimum. @Mathematics

OpenStudy (anonymous):

to express it in vertex form you need to make it look like this: f(x) = A*(x-h)^2+k so first factor out a 1/2... f(x) = 1/2(x^2+4x-12) now factor the parenthesis by completing the square. f(x) = 1/2*(x^2 + 4x - 12 +4 -4) f(x) = 1/2*(x+2)^2-8 This is vertex form.. where the vertex is (-2,-8) This is a local minimum because the graph opens up.